# DiP: Taming Diffusion Models in Pixel Space

### v1

@@ 2 Related Work
Latent Diffusion Models. Latent Diffusion Models (LDMs) [rombach2022high, podell2023sdxl, esser2024scaling, flux, xie2024sana, chen2023pixart, zhang2023adding] have become the de-facto paradigm for large-scale generative modeling due to their computational efficiency and scalability. By performing the diffusion process in a compressed latent space learned by a VAE, LDMs drastically reduce memory and computational costs, enabling efficient training and inference at high resolutions. Architectural advancements within this paradigm, such as replacing the U-Net [ronneberger2015u] backbone with a more scalable Transformer (DiT) [peebles2023scalable], have further pushed the boundaries of generation quality. Despite their success, this efficiency comes at a cost: the VAE acts as an information bottleneck, imposing a hard ceiling on the final image fidelity and often introducing subtle reconstruction artifacts [skorokhodov2025improving, yao2025reconstruction]. Our work circumvents these limitations by proposing an equally efficient architecture that operates directly in pixel space, thereby eliminating the VAE-induced quality constraints.
Pixel Diffusion Models. Recent years have seen renewed interest in pixel space diffusion models that aim to maximize signal fidelity while addressing computational inefficiency. Early works such as ADM [dhariwal2021diffusion] and DDPM [ho2020denoising] demonstrated the power of diffusion but were constrained by the quadratic complexity of their backbones, rendering them impractical for high resolutions. Multi-scale and image patch-based methods [hoogeboom2024simpler, ding2023patched, chen2025pixelflow, hoogeboom2023simple, hoogeboom2024simpler] further enhance the generation effect by decomposing large images into small image patches. However, these methods essentially simulate locality through brute-force training, which leads to extremely low efficiency. Recent work by PixelNerd [wang2025pixnerd] leverages a Transformer to process image features, which then conditions a NeRF-like coordinate network to act as a renderer for finely reconstructing each image patch, achieving impressive performance. Nevertheless, PixelNerd tightly couples the success of its method with this specific NeRF-like rendering mechanism, which may limit the exploration of a broader design space. We argue that the key to achieving efficient and high-quality pixel space generation lies not in the reliance on a specific structure like NeRF, but rather in the design principle of decoupling global structure construction from local detail refinement. Based on this insight, this paper aims to provide a more principled, efficient, and general solution for pixel space diffusion models.
@@ 3.3 Framework
DiP-XL/16 1.90 4.58 285.4 0.81 0.62 0.92s 320 100x2 631M
Table 1: Comparison of the performance of different methods on ImageNet 256×256 with CFG. Performance metrics are annotated with $\uparrow$ (higher is better) and $\downarrow$ (lower is better). Our method achieves the best FID score. Furthermore, compared to other pixel diffusion models, we achieve the best performance across all metrics with the lowest latency.
@@ 4.2 Main Result
Performance. Table 3.3 presents a comprehensive comparison against recent SOTA methods with classifier-free guidance scheduling with guidance interval [kynkaanniemi2024applying]. After 320 training epochs, DiP achieved an FID of 1.90 without requiring a pre-trained VAE, surpassing potentially diffusion models such as DiT-XL (FID 2.27) and SiT-XL (FID 2.06), which require longer training times. DiP outperforms the previous best pixel-based model, PixelFlow-XL/4 (FID 1.98), and significantly exceeds others like ADM (FID 3.94) and VDM++ (FID 2.12). Even with a shorter training schedule of 160 epochs, our model reaches a competitive FID of 2.16, outperforming established models like DiT-XL that require much longer training.

### v2

@@ 2 Related Work
Latent Diffusion Models. Latent Diffusion Models (LDMs) [rombach2022high, podell2023sdxl, esser2024scaling, flux, xie2024sana, chen2023pixart, zhang2023adding] have become the de-facto paradigm for large-scale generative modeling due to their computational efficiency and scalability. By performing the diffusion process in a compressed latent space learned by a VAE, LDMs drastically reduce memory and computational costs. Architectural advancements within this paradigm, such as replacing the U-Net [ronneberger2015u] backbone with a more scalable Transformer (DiT) [peebles2023scalable], have further pushed the boundaries of generation quality. Despite their success, this efficiency comes at a cost: the VAE acts as an information bottleneck, imposing a hard ceiling on the final image fidelity and often introducing subtle reconstruction artifacts [skorokhodov2025improving, yao2025reconstruction]. Our work circumvents these limitations by proposing an equally efficient architecture that operates directly in pixel space, thereby eliminating the VAE-induced quality constraints.
Pixel Diffusion Models. Recent years have seen renewed interest in pixel space diffusion models that aim to maximize signal fidelity while addressing computational inefficiency. Early works such as ADM [dhariwal2021diffusion] and DDPM [ho2020denoising] demonstrated the power of diffusion but were constrained by the quadratic complexity of their backbones, rendering them impractical for high resolutions. Multi-scale and image patch-based methods [hoogeboom2024simpler, ding2023patched, chen2025pixelflow, hoogeboom2023simple, hoogeboom2024simpler] further enhance the generation effect by decomposing large images into small patches. However, these methods essentially simulate locality through brute-force training, which leads to extremely low efficiency. Concurrent work JiT [li2025jit] demonstrates that high-dimensional data in pixel space can be effectively modeled by predicting clean images. Recent work by PixelNerd [wang2025pixnerd] leverages a Transformer to process image features, which then conditions a NeRF-like coordinate network to act as a renderer for finely reconstructing each image patch, achieving impressive performance. Nevertheless, PixelNerd tightly couples the success of its method with this specific NeRF-like rendering mechanism, which may limit the exploration of a broader design space. We argue that the key to achieving efficient and high-quality pixel space generation lies not in relying on a specific structure like NeRF, but rather in the design principle of decoupling global structure construction from local detail refinement. Based on this insight, this paper aims to provide a more principled, efficient, and general solution for pixel space diffusion models.
@@ 3.3 Framework
DiP-XL/16 1.79 4.59 281.9 0.80 0.63 0.92s 600 100x2 631M
Table 1: Comparison of the performance of different methods on ImageNet 256×256 with Euler solver and CFG. Performance metrics are annotated with $\uparrow$ (higher is better) and $\downarrow$ (lower is better). Our method achieves the best FID score. Furthermore, compared to other pixel diffusion models, we achieve the best performance across all metrics with the lowest latency.
@@ 4.2 Main Result
Performance. Table 3.3 presents a comprehensive comparison against recent SOTA methods with classifier-free guidance scheduling with guidance interval [kynkaanniemi2024applying]. After 600 training epochs, DiP achieved an FID of 1.79 without requiring a pre-trained VAE, surpassing potentially diffusion models such as DiT-XL (FID 2.27) and SiT-XL (FID 2.06), which require longer training times. DiP outperforms the previous best pixel-based model, PixelFlow-XL/4 (FID 1.98), and significantly exceeds others like ADM (FID 3.94) and VDM++ (FID 2.12). Even with a shorter training schedule of 160 epochs, our model reaches a competitive FID of 2.16, outperforming established models like DiT-XL that require much longer training.